# DART: Denoising Autoregressive Transformer for Scalable Text-to-Image Generation

### v1

@@ Abstract
Figure 1: Curated examples of images generated by DART at $256\times 256$ and $512\times 512$ pixels.
@@ 1 Introduction
In parallel, autoregressive models, such as GPT-4 (Achiam et al., 2023), have shown great success in modeling long-range dependencies in sequential data, particularly in the field of natural language processing. These models efficiently cache computations and manage dependencies across time steps, which has also inspired research into adapting autoregressive models for image generation. Early efforts such as PixelCNN (Oord et al., 2016) suffered from high computational costs due to pixel-wise generation. More recent models like VQ-GAN (Esser et al., 2021) and related work (Yu et al., 2022; Team, 2024; Tian et al., 2024) learn models of quantized images in a compressed latent space; Li et al. (2024) propose to generate directly in such space without quantization by employing a diffusion-based loss function. However, these methods fail to fully leverage the progressive denoising benefits of diffusion models, resulting in limited global context and error propagation during generation.
@@ 2.2 Autoregressive Models
In the field of natural language processing, Transformer models have achieved notable success in autoregressive modeling (Vaswani et al., 2017b; Raffel et al., 2020). Building on this success, similar approaches have been applied to image generation (Parmar et al., 2018; Esser et al., 2021; Chen et al., 2020; Yu et al., 2022; Sun et al., 2024; Team, 2024). Different from diffusion-based methods, these methods typically focus on learning the dependencies among discrete image tokens (e.g., through Vector Quantization (Van Den Oord et al., 2017)). To elaborate, consider an image ${\bm{x}}\in\mathbb{R}^{3\times H\times W}$. The process begins by encoding this image into a sequence of discrete tokens ${\bm{z}}_{1:N}=\mathcal{E}({\bm{x}})$. These tokens are designed to approximately reconstruct the original image through a learned decoder $\hat{{\bm{x}}}=\mathcal{D}({\bm{z}}_{1:N})$. An autoregressive model is then trained by maximizing the cross-entropy as follows:
@@ 3.3 Multi-resolution Generation
By this approach, the model can balance the number of noise levels with the total number of tokens to achieve better efficiency. Additionally, learning resolutions can be progressively increased by finetuning low-resolution models with extended sequences.
@@ 4.2 Results
Figure 7 also presents examples of generated results on ImageNet from all models using 16 sampling steps. DART-FM tends to produce sharper images with higher fidelity at higher CFG values. In contrast, DART-AR demonstrates an ability to generate more realistic samples at lower CFG values when compared to both the baselines and other variants. More samples are shown in Appendix C.
@@
To validate the scalability of high-resolution image generation, we further finetune DART by jointly modeling $256\times 256$ and $512\times 512$ resolutions within a single autoregressive process (§ 3.3). Sampled results are shown in Figure 6 and Appendix C.
@@ B.1 Architecture
In Matryoshka-DART upsample tuning, we follow the default setting and set the patch size as 2 for $512\times 512$ images, which end up containing 1024 tokens each. In total, the model has 8192 tokens in total, including 4096 tokens for $256\times 256$ images with 16 denoising steps and 4096 tokens for $512\times 512$ images with 4 denoising steps. We also applies rotary positional embedding (RoPE) (Su et al., 2024) to embed the information of token position. We find it critical to spatially align the positional embedding of high resolution images with low resolution ones. In particular, we define the rotary matrix feature at position $(i,j)$ in a $256\times 256$ image as $\mathbf{R}_{\Theta}^{(i,j)}$. For feature $x^{(i^{\prime},j^{\prime})}$ at position $(i^{\prime},j^{\prime})$ in the corresponding high resolution image with upsample ratio $r$, the rotary matrix is given $\mathbf{R}_{\Theta}^{(i^{\prime}/r,j^{\prime}/r)}$. In our case, the ratio is 2, which halves the position values on $512\times 512$ images.
@@ B.3 Parameterization of Diffusion Model
Appendix C Additional Samples
C.1 Prompts for Figure 1 (from left to right, top to bottom)
  • Cinematic photo of a fluffy baby Quokka with a knitted hat eating a large cup of popcorns, close up, studio lighting, screen reflecting in its eyes. 35mm photographs, film, bokeh, professional, 4k, highly detailed
  • A golden autumn scene, with the sun shining through a canopy of orange and red leaves, a path covered in fallen leaves, and soft light casting long shadows
  • Detailed pen and ink drawing of the arrival of a giant alien ship.
  • A diffusion cloud expanding in a cosmic scene, with the words ’DART’ subtly emerging from the nebula, blending into the stars and dust
  • A polar bear made entirely of vanilla ice cream, standing on a waffle cone iceberg with chocolate syrup dripping from its fur.
  • A lion made entirely out of swirling fire, standing majestically on a rocky outcrop as the sun sets behind it.
  • A pig made out of vegetables
  • At a futuristic wedding ceremony, a robot sharp and precise, holds the hand of the other robot.
  • A panda chef in a professional kitchen, wearing a chef’s hat and apron, carefully preparing a gourmet dish on the counter with steam rising from the pot.
  • In the image, a corgi dog is wearing a Santa hat and is laying on a fluffy rug. The dog’s tongue is sticking out and it appears to be happy. There are two Christmas boxes and a basket of toys nearby, indicating that the scene takes place during the winter season. The background features a Christmas tree, further suggesting the holiday atmosphere. The image has a warm and cozy feel to it, with the dog looking adorable in its hat and the boxes adding a festive touch
  • An astronaut standing on a frozen alien planet, looking up at a sky filled with two suns and a swirling aurora, with ice-covered
  • A lighthouse standing tall on a rocky shore, with massive waves crashing against the base, as the beam of light cuts through the night.
  • A scarecrow standing at the edge of a forest in early winter, its clothes dusted with fresh snow, with crows circling above against a cold, gray sky
  • The phrase ’GenAI’ made of scattered coffee beans that form the letters, arranged on top of a white kitchen table with a coffee mug.
  • An origami dragon made from rainbow-colored paper, soaring through a sky filled with clouds, leaving a trail of vibrant color behind it
  • Photography closeup portrait of an adorable rusty broken down steampunk robot covered in budding vegetation, surrounded by tall grass, misty futuristic sci-fi forest environment
@@ C.3 Additional Samples
We here show more text-to-image generative examples from DART-AR and DART-FM at resolution $256\times 256$ in Figure 13. We also show examples at resolution $512\times 512$ from Matryoshka-DART finetuned from DART-FM at $256\times 256$ in Figures 14 and 15.
@@
Figure 14: Uncurated samples from DART-FM with Matryoshka-DART fine-tuning on text-to-image generation at $512\times 512$ pixels given various captions.
Figure 15: Uncurated samples from DART-FM with Matryoshka-DART fine-tuning on text-to-image generation at $512\times 512$ pixels given various captions.

### v2

@@ Abstract
Figure 1: Curated examples of images generated by DART at $256^{2}$, $512^{2}$ and $1024^{2}$ pixels.
@@ 1 Introduction
In parallel, autoregressive models, such as GPT-4 (Achiam et al., 2023), have shown great success in modeling long-range dependencies in sequential data, particularly in the field of natural language processing. These models efficiently cache computations and manage dependencies across time steps, which has also inspired research into adapting autoregressive models for image generation. Early efforts such as PixelCNN (Oord et al., 2016) suffered from high computational costs due to pixel-wise generation. More recent models like VQ-GAN (Esser et al., 2021a) and related work (Yu et al., 2022; Team, 2024; Tian et al., 2024) learn models of quantized images in a compressed latent space; Li et al. (2024) propose to generate directly in such space without quantization by employing a diffusion-based loss function. However, these methods fail to fully leverage the progressive denoising benefits of diffusion models, resulting in limited global context and error propagation during generation.
@@ 2.2 Autoregressive Models
In the field of natural language processing, Transformer models have achieved notable success in autoregressive modeling (Vaswani et al., 2017b; Raffel et al., 2020). Building on this success, similar approaches have been applied to image generation (Parmar et al., 2018; Esser et al., 2021a; Chen et al., 2020; Yu et al., 2022; Sun et al., 2024; Team, 2024). Different from diffusion-based methods, these methods typically focus on learning the dependencies among discrete image tokens (e.g., through Vector Quantization (Van Den Oord et al., 2017)). To elaborate, consider an image ${\bm{x}}\in\mathbb{R}^{3\times H\times W}$. The process begins by encoding this image into a sequence of discrete tokens ${\bm{z}}_{1:N}=\mathcal{E}({\bm{x}})$. These tokens are designed to approximately reconstruct the original image through a learned decoder $\hat{{\bm{x}}}=\mathcal{D}({\bm{z}}_{1:N})$. An autoregressive model is then trained by maximizing the cross-entropy as follows:
@@ 3.3 Multi-resolution Generation
By this approach, the model can balance the the number of noise levels with the total number of tokens to achieve better efficiency. Additionally, learning resolutions can be progressively increased by finetuning low-resolution models with extended sequences.
@@ 4.2 Results
Table 1 lists the performance of our proposed DART in comparison to recent generative models on ImageNet-256. The reproduced results use the official codebase and checkpoint and we follow the best performing cfg scale reported in the original papers. We report reproduced results of DiT (Peebles & Xie, 2022), SiT (Ma et al., 2024), and MAR Li et al. (2024) with 16 sampling times which is the same as our vanilla DART. Our model achieves competitive performance when compared with diffusion models like LDM (Rombach et al., 2021) and AR models like VQGAN (Esser et al., 2021a) and RQ-Transformer (Lee et al., 2022). Admittedly, there is a gap between DART and SOTA visual generative models (like VAR (Tian et al., 2024) and MAR (Li et al., 2024)). However, we want to point out that many baselines are trained with significantly more FLOPs. For example, DiT (Peebles & Xie, 2022) is trained for 7M iterations whereas DART is only trained for 500k iterations. Also, baselines like VAR and MAR employ larger models than our DART. In particular, VAR deploys a 2B model while the largest DART model is approximately 800M. Besides, in our reproduced results, when using only 16 sampling steps as the setting of our vanilla DART, our model show significantly better performance than DiT, SiT and MAR. Also, we report MAR-AR (Li et al., 2024), a variant of MAR which generate tokens in an autoregressive manner instead of masked modeling which is applied in standard MAR models. DART which generates samples through autoregressive denoising shows better performance than MAR-AR. These results further validate the effectiveness of leveraging the whole denoising trajectory.
Table 1: Generative models on class-conditional ImageNet $256\times 266$. *: reproduced from official codebase and checkpoints.
[…]
Diff. ADM (Dhariwal & Nichol, 2021) 10.94 101.0 554M 250
CDM (Ho et al., 2022) 4.88 158.7 - 8100
LDM (Rombach et al., 2021) 3.60 247.7 400M 250
DiT (Peebles & Xie, 2022) 2.27 278.2 675M 250
SiT (Ma et al., 2024) 2.06 277.5 675M 250
AR VQGAN (Esser et al., 2021b) 15.78 74.3 1.4B 256
RQTran (Lee et al., 2022) 3.80 323.7 3.8B 68
MAR-AR (Li et al., 2024) 4.69 244.6 479M 256
MAR (Li et al., 2024) 1.55 303.7 943M 256
VAR (Tian et al., 2024) 1.73 350.2 2.0B 10
Reprod. DiT* (Peebles & Xie, 2022) 19.52 125.9 675M 16
SiT* (Ma et al., 2024) 6.98 122.9 675M 16
MAR* (Li et al., 2024) 6.37 221.3 943M 16
Ours DART 5.62 231.7 812M 16
DART-AR 3.98 256.8 812M 4096
DART-FM 3.82 263.8 820M 16
Figure 7 also presents examples of generated results on ImageNet from all models using 16 sampling steps. DART-FM tends to produce sharper images with higher fidelity at higher CFG values. In contrast, DART-AR demonstrates an ability to generate more realistic samples at lower CFG values when compared to both the baselines and other variants. More samples are shown in Appendix D.
@@
To validate the scalability of high-resolution image generation, we further finetune DART by jointly modeling $256\times 256$ and $512\times 512$ resolutions within a single autoregressive process (§ 3.3). Sampled results are shown in Figure 6 and Appendix D.
@@ B.1 Architecture
In Matryoshka-DART upsample tuning, we follow the default setting and set the patch size as 2 for $512\times 512$ images, which end up containing 1024 tokens each. In total, the model has 8192 tokens in total, including 4096 tokens for $256\times 256$ images with 16 denoising steps and 4096 tokens for $512\times 512$ images with 4 denoising steps. We also applies rotary positional embedding (RoPE) (Su et al., 2024) to embed the information of token position. We find it critical to spatially align the positional embedding of high resolution images with low resolution ones. In particular, we define the rotary matrix for feature $x^{(i,j)}$ at position $(i,j)$ in a $256\times 256$ image as $\mathbf{R}_{\Theta}^{(i,j)}$. For feature $x^{(i^{\prime},j^{\prime})}$ at position $(i^{\prime},j^{\prime})$ in the corresponding high resolution image with upsample ratio $r$, the rotary matrix is given $\mathbf{R}_{\Theta}^{(i^{\prime}/r,j^{\prime}/r)}$. In our case, the ratio is 2, which halves the position values on $512\times 512$ images.
@@ B.3 Parameterization of Diffusion Model
Appendix C Comparison of DART Variants
We here further clarify the differences and connections between the variants of DART. Table 3 lists the major comparison between DART, DART-AR, and DART-FM. Conceptually, DART predicts the denoised value and adds independent noise to acquire a less noisy image at each step. It conducts this denoising process autoregressively until the clean image is generated. DART-AR applies a token-wise autoregression instead of block-wise autoregressive in vanilla DART, which conduct denoising generation in a more fine-grained granularity. DART-FM, on the other hand, keeps the block-wise autoregression while introduces an additional flow network to conduct flow-mating-based refinement for generated tokens. Both DART-AR and DART-FM improve the performance over vanilla DART. In general, DART-FM demonstrates a better tradeoff between generation quality and inference efficiency. Table 3 summarizes the difference between DART variants.
Kaleido-DART is for multimodal text-image generation, which integrates next-token prediction for text and next-denoising prediction for images (proposed in our DART). In image generation, it can seamlessly adapt all three variants of proposed methods: DART, DART-AR, DART-FM. Similarly, Matryoshka-DART, which enables multi-resolution generation, also adapts to all the three DART variants. Since in Matryoshka-DART, one can simply concatenate high-resolution image tokens after the low-resolution ones, which doesn’t affect the denoising modeling in these variants.
Table 3: Comparison of DART variants.
Model Attn Mask #AR steps #FM steps
DART Block-wise Causal 16 0
DART-AR Causal 4096 0
DART-FM Block-wise Causal 16 1600
Appendix D Additional Samples
D.1 Step-wise Generative Results
@@ D.2 Additional Samples
We here show more text-to-image generative examples from DART-AR and DART-FM at resolution $256\times 256$ in Figure 13. We also show examples at resolutions $512\times 512$ and $1024$ from Matryoshka-DART finetuned from DART-FM at $256\times 256$ in Figures 14 and 15.
@@
Figure 14: Uncurated samples from DART-FM with Matryoshka-DART fine-tuning on text-to-image generation at $512\times 512$ and $1024$ pixels given various captions.
Figure 15: Uncurated samples from DART-FM with Matryoshka-DART fine-tuning on text-to-image generation at $512\times 512$ and $1024$ pixels given various captions.